Write the metric unit that makes each statement true.
(1.) 850 g = 0.85
(2.) 17.43 m = 1,743
(3.) 12,451 dL = 1,245.1
(4.) 2.48 cm = 24.8
thank you!

Answer by stanbon(48564)

Write the metric unit that makes each statement true.
(1.) 850 g = 0.85*1000g = 0.85 kg
(2.) 17.43 m = 1,743 *(1/100)m = 1743 cm
(3.) 12,451 dL = 1,245.1 ?
Conversion tables say there is no such thing as dL.
(4.) 2.48 cm = 24.8 (1/10)cm = 24.8 mm
===========
Cheers,
Stan H.